∫ ∫ π‘₯𝑦𝑑π‘₯𝑑𝑦 2𝑦


𝑦


2





1
Expert's answer
2021-12-09T16:03:13-0500

In unit circle we need to find the integral value.

"\int \int xydxdy= \int_0^{2\pi}\int_0^1r^3sin\theta cos\theta drd\theta\\
=\int_0^{2\pi}[\frac{r^4}{4}]_0^1sin\theta cos\theta drd\theta\\
=\int_0^{2\pi}\frac{1}{4}sin\theta cos\theta drd\theta\\
=\int_0^{2\pi}\frac{1}{8}2sin\theta cos\theta drd\theta\\
=\int_0^{2\pi}\frac{1}{8}sin(2\theta)drd\theta\\
=\frac{1}{8}[\frac{-cos(2\theta)}{2}]_0^{2\pi}\\
=\frac{1}{16}[-1+1]\\
=0"